# Ratio Covers of Convex Sets and Optimal Mixture Density Estimation

### v1

@@ 4.2 Approximate Pareto curves
$|A|\leqslant\left(c\,\frac{\log R}{\varepsilon}\right)^{d-1}$
@@
By comparison, we now use our ratio cover theorem to obtain an $\varepsilon$-approximate Pareto curve whose size has no dependence on $R$ and whose base does not involve $d$. The convexity assumption is essential for this stronger statement (see the discussion following Theorem˜1), and it is natural in many applications, for example when $K$ is the feasible region of a linear program.
Corollary 2 (of Theorem˜1).
@@
In particular, this implies $|A|\leqslant\left(\frac{c_{1}}{\varepsilon}\right)^{c_{2}d}$ for large enough absolute constants $c_{1},c_{2}>0$. The proof of Corollary˜2 is deferred to Appendix˜D. A substantial literature following Papadimitriou and Yannakakis develops algorithms that construct $\varepsilon$-approximate Pareto curves of near-minimal size (see, e.g., Vassilvitskii and Yannakakis (2005); Diakonikolas and Yannakakis (2008, 2010); Bazgan et al. (2015); Daskalakis et al. (2016)). Our result shows that, under convexity, there always exists an $\varepsilon$-approximate Pareto curve of smaller size, and therefore yields improved existential upper bounds for what such approximation algorithms can aim to achieve in this setting.
4.3 Proof of a version of Theorem˜1 based on the reduction to $\varepsilon$-nets
In this section, we provide an alternative proof of Theorem˜1 based on $\varepsilon$-net arguments in the spirit of Haussler and Welzl Haussler and Welzl (1986), introduced in the context of Pareto curves by Diakonikolas and Yannakakis Diakonikolas and Yannakakis (2010) for general sets $K$. We then combine this approach with convexity via a volumetric argument similar to the one used in the proof of Theorem˜3. We state the resulting bound as a separate theorem because it yields a different dependence on $\varepsilon$ in the language of Corollary˜2.
@@ 4.3 Proof of a version of Theorem˜1 based on the reduction to $\varepsilon$-nets
For every convex and compact set $K\subset\mathbb{R}_{+}^{d}$ with nonempty interior and every $0<\varepsilon\leqslant 1$, there exists an $\varepsilon$-approximate Pareto curve $A\subset K$ with
@@
Note that, in full generality, the dependence on $\varepsilon$ in Theorem˜5 differs from that in Corollary˜2. In the natural regime of fixed $d$ and sufficiently small $\varepsilon$, the bound of Corollary˜2 is stronger because it has exponent d1d-1 rather than $d$. Despite not having the same dependence on $\varepsilon$ as Corollary˜2, the result of Theorem˜5 implies a $32$-ratio cover of size at most
@@
for all $d\geqslant 1$. Since optimization of the numerical constants is not our priority, we mostly use the simple integer form of Theorem˜1.
We also note that the construction in our first proof of Corollary˜2 is deterministic and constructive, whereas the proof of Theorem˜5 uses a Vapnik–Chervonenkis style probabilistic argument.
@@
Thus, it suffices to construct a set $A$ such that for all $\theta\in K$ one has $A\cap R_{\theta}(\varepsilon)\neq\emptyset$, and such that every range satisfies the uniform lower bound $\mu(R_{\theta}(\varepsilon))\geqslant\eta$, where $\eta=\left(\frac{\varepsilon}{1+\varepsilon}\right)^{d}.$ By the standard $\varepsilon$-net argument (see Haussler and Welzl (1986) and (Blumer et al., 1989, Theorem 2.1)), if we sample
@@ 5 Mixture models (convex aggregation)
  • Use the first $n$ observations to run a Birgé-Le Cam tournament (see Theorem 9) over $\mathcal{P}$ to find a convex closed set $\mathcal{W}\subset\mathcal{P}$ such that, with probability $1-\delta/2$, it satisfies
    $\sup\limits_{p,q\in\mathcal{W}}\operatorname{H}^{2}(p,q)\leqslant\frac{3528\log(64)\,M+512\log(4/\delta)}{n},\quad\textrm{and}\quad p^{\star}\in\mathcal{W}.$ (9)
  • Let $\Delta$ be the simplex in $\mathbb{R}^{M}$. Consider the closed convex set $\Delta^{\prime}\subseteq\Delta$ defined as follows:
    $\Delta^{\prime}=\left\{\alpha\in\Delta:\sum_{j=1}^{M}\alpha_{j}p_{j}\in\mathcal{W}\right\},$
    and construct a $32$-ratio cover of $\Delta^{\prime}$, which will be denoted as $\Delta^{\prime}_{c}$.
  • Choose any $\beta^{(1)},\ldots,\beta^{(M)}\in\Delta^{\prime}$ such that for all $j\in[M]$, $\beta^{(j)}_{j}=\sup\{\alpha_{j}:\alpha\in\Delta^{\prime}\}$, i.e., each vector $\beta^{(j)}$ witnesses the largest $j$-th coordinate in $\Delta^{\prime}$. Construct a (random) density class
    $\mathcal{W}^{\prime}=\left\{\frac{1}{2}\sum\limits_{j=1}^{M}\alpha_{j}p_{j}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}:\alpha\in\Delta^{\prime}_{c}\right\}.$ (10)
  • Using the second half of the sample run a version of the progressive density estimator of Lemma˜2 over $\mathcal{W}^{\prime}$ to output the final density $\widehat{p}$.
@@
Fix $\delta\in(0,1)$. Combining the Birgé-Le Cam tournament result (Theorem 9) and the local Hellinger entropy bound of Theorem 3, we construct the desired convex set $\mathcal{W}\subset\mathcal{P}$ satisfying, with probability at least $1-\delta/2$,
$\sup\limits_{p,q\in\mathcal{W}}\operatorname{H}^{2}(p,q)\leqslant\frac{3528\log(64)\,M+512\log(4/\delta)}{n},\quad\textrm{and}\quad p^{\star}\in\mathcal{W}.$ (11)
From now on we work on this event and focus on the second half of the sample. Using the second half of the sample we output the density $\widehat{p}$ obtained using the procedure of Lemma˜2 with respect to $\mathcal{W}^{\prime}$ defined by (10). Let $p^{\star}=\sum\limits_{j=1}^{M}\alpha_{j}p_{j}$ for some $\alpha\in\Delta^{\prime}$. On the event where $p^{\star}\in\mathcal{W}$, by the definition of $\Delta^{\prime}_{c}$ there is a vector $\gamma\in\Delta^{\prime}_{c}$ such that for the density
$p^{\star}_{c}=\sum\limits_{j=1}^{M}\gamma_{j}p_{j},\quad\textrm{it holds that}\quad\max_{j,\gamma_{j}>0}\frac{\alpha_{j}}{\gamma_{j}}\leqslant 32.$ (12)
@@
$\sup_{p_{\alpha},p_{\alpha^{\prime}}\in\mathcal{W}^{\prime}}\sup_{x\in\mathcal{X}}\log\left(\frac{p_{\alpha}(x)}{p_{\alpha^{\prime}}(x)}\right)\leqslant\log\left(1+\frac{\sum_{k}\alpha_{k}p_{k}(x)}{\sum_{k}\bar{\beta}_{k}p_{k}(x)}\right)\leqslant\log\left(1+M\right)\leqslant\log(2M).$
Therefore, by Lemma˜2, and since by Theorem 1 it holds that $|\mathcal{W}^{\prime}|\leqslant 256^{M}$, we have, with probability at least $1-\delta/2$,
@@
Finally, by Lemma 1 and the convexity of the Hellinger distance squared, and observing that the ratio of $p^{\star}$ to $\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}$ is bounded by $64$ by (12), we note explicitly that since $\alpha_{j}\leqslant 32\,\gamma_{j}$ coordinatewise,
$\sum_{j=1}^{M}\alpha_{j}p_{j}(x)\leqslant 32\sum_{j=1}^{M}\gamma_{j}p_{j}(x)=32\,p_{c}^{\star}(x),\quad\text{so}\quad\frac{p^{\star}(x)}{\widetilde{p}(x)}\leqslant\frac{32p_{c}^{\star}(x)}{\frac{1}{2}p_{c}^{\star}(x)}=64.$
Therefore,
@@
where in the last line we used that $p^{\star}\in\mathcal{W}$ and $\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}\in\mathcal{W}$ combined with (11). Plugging this into (13) and using the union bound, we have with probability at least $1-\delta$,
@@
Note that we used a sample of size $2n$, which leads to an increased absolute constant by at most a factor of $3$ (e.g., in the worst case of an odd total sample size such as $3$ one observation is dropped). The claim follows.∎
@@ 6 Applications of ratio covers
6.1 Sharpening the Yang–Barron bound (expected error guarantees)
@@
Define the progressive-mixture estimator $\widehat{p}=\sum_{j=1}^{N}w_{j}q_{j}$, where
@@
Split the sample into two independent halves. Let $\widehat{p}_{\operatorname{H}}$ be the Le Cam–Birgé estimator (Theorem 9) built from the first $n/2$ observations, and let $\widehat{p}$ be the Yang–Barron mixture (Theorem 7) built from the remaining $n/2$ observations over the candidate set $S(\widehat{p}_{\operatorname{H}})$ with the uniform prior. Conditioning on the first half and applying Theorem 7 to the second half yields
@@
For $x\in(0,1)$, Theorem 9 applied with sample size $n/2$ gives, with probability at least 1x1-x,
@@
This immediately implies the following sharp result based on the progressive mixture rule of Yang and Barron applied to a localized set of densities.
@@
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant\frac{10568316\,M}{n}.$
@@ 6.2 KL covering numbers
$\chi^{2}$ method $0<\varepsilon<\dfrac{1}{4(d+1)^{2}}$ $\left(\dfrac{c}{\varepsilon}\right)^{\frac{d-1}{2}}$ $\bigl(\log\frac{1}{\varepsilon}\bigr)\left(\dfrac{C}{\varepsilon}\right)^{\frac{d-1}{2}}$
@@
Fix $\varepsilon>0$ and set $R=\varepsilon/70$. Let $\{q^{(1)},\dots,q^{(N)}\}\subset\Delta_{d-1}$ be an $R$-cover of $\Delta_{d-1}$ in squared Hellinger distance, so that for every $p\in\Delta_{d-1}$ there exists $k\in[N]$ with $\operatorname{H}^{2}(p,q^{(k)})\leqslant R$, and $N=N_{\operatorname{H}^{2}}(\Delta_{d-1},R)$. For each $k\in[N]$, define the (squared-Hellinger) ball intersection
@@
Applying Theorem˜1 to a convex and compact set $K_{k}$, we obtain a $32$-ratio cover $A_{k}\subset K_{k}$ with $|A_{k}|\leqslant 2^{8d}$. Define $\mathcal{A}=\bigcup_{k=1}^{N}A_{k}$, so
@@ 7 Aggregation in the misspecified case
  • the analysis relies on more standard tools;
  • the resulting algorithms are simpler;
@@
  • at the price of additional polylogarithmic factors in $n$ and $M$ in the bounds.
Thus, in the broader context of our results, the hardest remaining question appears to be obtaining optimal rates up to multiplicative constant factors when $p^{\star}\notin\mathcal{P}$. Regarding algorithmic simplicity, in the context of convex aggregation we will reduce the problem to a classical MLE analysis over the convex class, which might have computational advantages for particular families of densities.
@@
Compared to Theorem˜2 the bound contains an additional dependence on $n$ under the logarithm. The estimator and the analysis are presented in Appendix˜D. The idea behind the estimator is to run the predictor of Lemma˜2 on a deterministically smoothed class $\mathcal{W}$ obtained from the given set $\mathcal{P}$.
@@
The main difference from Theorem˜6 is an additional squared logarithmic term in the risk bound. Since the estimator in Proposition˜4 is particularly simple, we present it.
@@
    $\mathcal{W}=\bigg\{\left(1-\frac{1}{n}\right)p+\frac{1}{nM}\sum_{q\in\mathcal{P}}q:\ p\in\operatorname{conv}(\mathcal{P})\bigg\}.$
  • Let $\widehat{p}$ be the MLE in $\mathcal{W}$. That is,
    $\widehat{p}=\arg\min_{p\in\mathcal{W}}\widehat{L}_{n}(p).$
The proof of Proposition˜4 follows relatively standard lines using the results in empirical process theory and we defer it to Appendix˜D.

### v2

@@ 4.2 Approximate Pareto curves
$|A|\leqslant\max\left\{1,\left(c\,\frac{\log R}{\varepsilon}\right)^{d-1}\right\}$
@@
By comparison, we now use our ratio cover theorem to obtain an $\varepsilon$-approximate Pareto curve whose size has no dependence on $R$ and whose base does not involve $d$. The convexity assumption is essential for this stronger statement (see the discussion following Theorem˜3), and it is natural in many applications, for example when $K$ is the feasible region of a linear program.
Corollary 2 (of Theorem˜3).
@@
The proof of Corollary˜2 is deferred to Appendix˜D. A substantial literature following Papadimitriou and Yannakakis develops algorithms that construct $\varepsilon$-approximate Pareto curves of near-minimal size (see, e.g., Vassilvitskii and Yannakakis (2005); Diakonikolas and Yannakakis (2008, 2010); Bazgan et al. (2015); Daskalakis et al. (2016)). Our result shows that under convexity, there always exists an $\varepsilon$-approximate Pareto curve of smaller size, and therefore yields improved existential upper bounds for what such approximation algorithms can aim to achieve in this setting.
4.3 Proof of a version of Theorem˜3 based on the reduction to $\varepsilon$-nets
In this section, we provide an alternative proof of Theorem˜3 based on $\varepsilon$-net arguments in the spirit of Haussler and Welzl Haussler and Welzl (1986), introduced in the context of Pareto curves by Diakonikolas and Yannakakis Diakonikolas and Yannakakis (2010) for general sets $K$. We then combine this approach with convexity via a volumetric argument similar to the one used in the proof of Theorem˜2. We state the resulting bound as a separate theorem because it yields a different dependence on $\varepsilon$ in the language of Corollary˜2.
@@ 4.3 Proof of a version of Theorem˜3 based on the reduction to $\varepsilon$-nets
For every convex and compact set $K\subset\mathbb{R}_{+}^{d}$ with nonempty interior and every $\varepsilon>0$, there exists an $\varepsilon$-approximate Pareto curve $A\subset K$ with
@@
Note that, in full generality, the dependence on $\varepsilon$ in Theorem˜5 differs from that in Corollary˜2. For instance, in the regime of fixed $d$ and sufficiently small $\varepsilon$, the bound of Corollary˜2 is stronger because it has exponent d1d-1 rather than $d$. Despite not having the same dependence on $\varepsilon$ as Corollary˜2, the result of Theorem˜5 implies a $32$-ratio cover of size at most
@@
for all $d\geqslant 1$. Since optimization of numerical constants is not our priority, we mostly use the simple integer form of Theorem˜3.
We also note that the construction in our first proof of Corollary˜2 is deterministic and constructive, whereas the proof of Theorem˜5 is nonconstructive and based on a probabilistic argument.
@@
Thus, it suffices to construct a set $A$ such that for all $\theta\in K$ one has $A\cap R_{\theta}(\varepsilon)\neq\emptyset$. Now the class $\mathcal{R}(\varepsilon)$ has VC dimension at most $d$, and every range satisfies the uniform lower bound $\mu(R_{\theta}(\varepsilon))\geqslant\eta$, where $\eta=\left(\frac{\varepsilon}{1+\varepsilon}\right)^{d}.$ Thus by a standard VC lower bound (see Haussler and Welzl (1986) and (Blumer et al., 1989, Theorem 2.1)), if we sample
@@ 5 Mixture models (convex aggregation)
  • Use the first $n$ observations to run a Birgé-Le Cam tournament (see Lemma˜9) over $\mathcal{P}$ to find a convex closed set $\widehat{\mathcal{Q}}_{1}\subset\mathcal{P}$ such that, with probability $1-\delta/2$, it satisfies
    $\sup\limits_{p,q\in\widehat{\mathcal{Q}}_{1}}\operatorname{H}^{2}(p,q)\leqslant\frac{3528\log(64)\,M+512\log(4/\delta)}{n},\quad\textrm{and}\quad p^{\star}\in\widehat{\mathcal{Q}}_{1}.$ (13)
  • Let $\Delta$ be the simplex in $\mathbb{R}^{M}$. Consider the closed convex set $\widehat{\Delta}\subseteq\Delta$ defined as follows:
    $\widehat{\Delta}=\left\{\theta\in\Delta:\sum_{j=1}^{M}\theta_{j}p_{j}\in\widehat{\mathcal{Q}}_{1}\right\},$
    and construct a $32$-ratio cover of $\widehat{\Delta}$ with cardinality at most $2^{8M}$, denoted as $\widehat{\Delta}_{c}$.
  • Choose any $\beta^{(1)},\ldots,\beta^{(M)}\in\widehat{\Delta}$ such that for all $j\in[M]$, $\beta^{(j)}_{j}=\sup\{\theta_{j}:\theta\in\widehat{\Delta}\}$, i.e., each vector $\beta^{(j)}$ maximizes the $j$-th coordinate in $\widehat{\Delta}$. Construct a (random) density class
    $\widehat{\mathcal{Q}}_{2}=\left\{\frac{1}{2}\sum\limits_{j=1}^{M}\theta_{j}p_{j}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}:\theta\in\widehat{\Delta}_{c}\right\}.$ (14)
  • Using the second half of the sample, run the density estimator of Lemma˜2 over $\widehat{\mathcal{Q}}_{2}$ to output the final density $\widehat{p}$.
@@
Fix $\delta\in(0,1)$. Combining the Birgé-Le Cam tournament result (Lemma˜9) and the local Hellinger entropy bound of Theorem˜2, we construct the desired convex set $\widehat{\mathcal{Q}}_{1}\subset\mathcal{P}$ satisfying, with probability at least $1-\delta/2$,
$\sup\limits_{p,q\in\widehat{\mathcal{Q}}_{1}}\operatorname{H}^{2}(p,q)\leqslant\frac{3528\log(64)\,M+512\log(4/\delta)}{n},\quad\textrm{and}\quad p^{\star}\in\widehat{\mathcal{Q}}_{1}.$ (15)
From now on we work under this event.
Using the second half of the sample, we output the density $\widehat{p}$ obtained using the procedure of Lemma˜2 with respect to $\widehat{\mathcal{Q}}_{2}$ defined by (14). Let $\theta\in\widehat{\Delta}$ be such that $p^{\star}=\sum\limits_{j=1}^{M}\theta_{j}p_{j}$. By definition of $\widehat{\Delta}_{c}$, there exists a vector $\phi\in\widehat{\Delta}_{c}$ such that
$\theta_{j}\leqslant 32\,\phi_{j}\quad\text{for every }j=1,\dots,M.$ (16)
We then define
$p^{\star}_{c}=\sum_{j=1}^{M}\phi_{j}p_{j}\quad\text{and}\quad\widetilde{p}=\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}.$
@@
$\sup_{p_{\alpha},p_{\alpha^{\prime}}\in\widehat{\mathcal{Q}}_{2}}\sup_{x\in\mathcal{X}}\log\left(\frac{p_{\alpha}(x)}{p_{\alpha^{\prime}}(x)}\right)\leqslant\log\left(1+\frac{\sum_{k}\alpha_{k}p_{k}(x)}{\sum_{k}\bar{\beta}_{k}p_{k}(x)}\right)\leqslant\log\left(1+M\right)\leqslant\log(2M).$
Therefore, by Lemma˜2, and since by Theorem 3 it holds that $|\widehat{\mathcal{Q}}_{2}|\leqslant 256^{M}$, we have, with probability at least $1-\delta/2$,
@@
Finally, by Lemma 1 and the convexity of the Hellinger distance squared, and observing that the ratio of $p^{\star}$ to $\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}$ is bounded by $64$ by (16), and, in particular, $p^{\star}(x)\leqslant 32\,p^{\star}_{c}(x).$ Therefore,
@@
where in the last line we used that $p^{\star}\in\widehat{\mathcal{Q}}_{1}$ and $\frac{1}{2}p^{\star}_{c}+\frac{1}{2M}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{M}\beta^{(j)}_{k}p_{k}\in\widehat{\mathcal{Q}}_{1}$ combined with (15). Plugging this into (17) and using the union bound, we have with probability at least $1-\delta$,
@@
Note that we used a sample of size $2n$, which leads to an increased absolute constant by at most a factor of $3$. The claim follows.∎
@@ 6 Applications of ratio covers
6.1 Expected error guarantees: sharpening the Yang–Barron bound
@@
Define the progressive mixture estimator $\widehat{p}=\sum_{j=1}^{N}w_{j}q_{j}$, where
@@
Split the sample into two independent halves. Let $\widehat{p}_{\operatorname{H}}$ be the Le Cam–Birgé estimator (Lemma˜9) built from the first $n/2$ observations, and let $\widehat{p}$ be the Yang–Barron mixture (Theorem 7) built from the remaining $n/2$ observations over the candidate set $S(\widehat{p}_{\operatorname{H}})$. Conditioning on the first half and applying Theorem 7 to the second half yields
@@
For $x\in(0,1)$, Lemma˜9 applied with sample size $n/2$ gives, with probability at least 1x1-x,
@@
When combined with the covering estimates of Theorem˜2 and Theorem˜3, this establishes the optimal in-expectation rate for mixture density estimation:
@@
${\mathbf{E}}\big[\operatorname{KL}(p^{\star},\widehat{p})\big]\leqslant 11000000\,\frac{M}{n}.$
@@ 6.2 KL covering numbers
$\chi^{2}$ method $0<\varepsilon<\dfrac{1}{4(d+1)^{2}}$ $\left(\dfrac{c}{\varepsilon}\right)^{\frac{d-1}{2}}$ $\log\left(\dfrac{1}{\varepsilon}\right)\left(\dfrac{C}{\varepsilon}\right)^{\frac{d-1}{2}}$
@@
If $d=1$, then the set is the singleton and the proof is immediate. Thus, we assume $d\geqslant 2$. Fix $\varepsilon>0$ and set $R=\varepsilon/70$. Let $\{q^{(1)},\dots,q^{(N)}\}\subset\Delta_{d-1}$ be an $R$-cover of $\Delta_{d-1}$ in squared Hellinger distance, so that for every $p\in\Delta_{d-1}$ there exists $k\in[N]$ with $\operatorname{H}^{2}(p,q^{(k)})\leqslant R$, and $N=N_{\operatorname{H}^{2}}(\Delta_{d-1},R)$. For each $k\in[N]$, define the (squared-Hellinger) ball intersection
@@
Applying Theorem˜3 to a convex and compact set $K_{k}$, we obtain a $32$-ratio cover $A_{k}\subset K_{k}$ with $|A_{k}|\leqslant 2^{8d}$. Define $\mathcal{A}=\bigcup_{k=1}^{N}A_{k}$, so
@@ 7 Aggregation in the misspecified case
  • the algorithms are simpler, and their analysis relies on less sophisticated arguments;
  • however, their statistical guarantees feature additional polylogarithmic factors in $n$ and $M$.
In particular, obtaining optimal rates up to universal constant factors when $p^{\star}\notin\mathcal{P}$ remains an open question.
@@
Compared to Theorem˜1 the bound contains an additional dependence on $n$ under the logarithm. The estimator and the analysis are presented in Appendix˜D. The idea behind the estimator is to run the predictor of Lemma˜2 on a deterministically smoothed class $\mathcal{Q}$ obtained from the given set $\mathcal{P}$.
@@
The main difference with Theorem˜6 consists in additional squared logarithmic terms in the risk bound. The estimator in Proposition˜4 corresponds to the MLE over a smoothed density class:
@@
    $\mathcal{Q}=\bigg\{\left(1-\frac{1}{n}\right)p+\frac{1}{nM}\sum_{q\in\mathcal{P}}q:\ p\in\operatorname{conv}(\mathcal{P})\bigg\}.$
  • Let $\widehat{p}$ be the MLE in $\mathcal{Q}$. That is,
    $\widehat{p}=\arg\min_{p\in\mathcal{Q}}\widehat{L}(p).$
The proof of Proposition˜4 follows relatively standard lines using techniques from empirical process theory and we defer it to Appendix˜D.